For the two sample Ztest of testing difference of means, the value of test statistics is z=1.25 and the test is two tailed. The p-value for the above test is

1
Expert's answer
2022-01-04T11:29:24-0500

"P(|Z|>1.25) = 2(1-P(Z<1.25)) \\

= 2 (1-0.8943) \\

= 0.2114"

The EXCEL formula to find the p-value for a two-tailed z-test is “=2*(1-NORM.S.DIST(1.25, TRUE))”.

=0.2114